22.2 g 33.3 g g g Limiting reactant:
2 g 15.0 g COCT
Answer #1

PART 1 : ANSWER : MASS OF NbC = 26.23 g and MASS OF CO = 17.5 g

Consider a reaction : 14 C + 2 Nb2O5  media%2Fdf5%2Fdf526095-f49b-41fc-a81a-9d 4 NbC + 10 CO

From reaction , stoichiometric ratio of reactants is C : Nb2O5   = 14: 2= 7: 1

No of moles of C = mass / Molar mass

= 22.2 g / 12.01 g / mol

= 1.85 mol

Molar mass of Nb2O5 = ( 2 x 92.91 ) + ( 5 x 16.00 ) = 265.8 g / mol

No of moles of Nb2O5 = 33.3 g / 265.8 g / mol

= 0.125 mol

Provided molar ratio of reactants is C : Nb2O5   = 1.85 : 0.125 = 14.8 : 1

Comparing provided molar ratio with stoichiometric ratio , it is found that Nb2O5 is limiting reactant and yield of product depends upon mass of Nb2O5.

From reaction , 2 moles Nb2O5  \equiv 4 mole NbC

Therefore, 0.125 moles Nb2O5  \equiv 4 x 0.125 / 2  mole NbC

\equiv 0.25 mole

Molar mass of NbC= 92.91 + 12.01 = 104.9 g / mol

0.25 mol NbC = 0.25 mol x ( 104.9 g / 1 mol )

= 26.23 g

From reaction , 2 moles Nb2O5  \equiv 10 mole CO

Therefore, 0.125 moles Nb2O5  \equiv 10 x 0.125 / 2  mole CO

\equiv 0.625 mole CO

Molar mass of CO =12.01 + 16.00 = 28.01 g / mol

0.625 mol CO = 0.625 mol x ( 28.01 g / 1 mol )

= 17.5 g

PART 2 :

Consider reaction : 3 Cl 2 + 6 NaOH media%2Fdf5%2Fdf526095-f49b-41fc-a81a-9d 5 NaCl + NaClO 3  + 3 H2O

Moles of NaOH = Mass / Molar mass

= 15.0 g / 40.0 g/ mol

= 0.35 mol

From reaction , 6 moles NaOH  \equiv3 mole Cl 2

Therefore, 0.35 moles NaoH  \equiv 3 x 0.35 / 6 mole Cl 2

\equiv 0.175 mole Cl 2

Molar mass of Cl 2 = 2 x 35.45 = 70.9 g / mol

0.175 mole Cl 2 = 0.175 mol x ( 70.9 g / 1 mol ) = 12.41 g

From reaction , 6 moles NaOH  \equiv5 mole NaCl

Therefore, 0.35 moles NaoH  \equiv 5 x 0.35 / 6 mole NaCl

\equiv 0.292 mole NaCl

Molar mass of NaCl = 22.99+ 35.45 = 58.44 g / mol

0.292 mole NaCl = 0.292 mol x ( 58.44 g / 1 mol ) =17.06 g

From reaction , 6 moles NaOH  \equiv1 mole NaClO3

Therefore, 0.35 moles NaoH  \equiv 1 x 0.35 / 6 mole NaClO3

\equiv 0.0583 mole NaClO3

Molar mass of NaClO3 = 22.99+ 35.45 + ( 3 x 16.00) = 106.44 g / mol

0.0583 mole NaClO3 = 0.0583 mol x ( 106.44 g / 1 mol ) =6.20 g

From reaction , 6 moles NaOH  \equiv3 mole H2O

Therefore, 0.35 moles NaoH  \equiv3 x 0.35 / 6 mole H2O

\equiv0.175 mole H2O

Molar mass of H2O = ( 2 x 1.0079 ) + 16.00 = 18.02 g / mol

0.175 mole H2O = 0.175 mol x ( 18.02 g / 1 mol ) = 3.15 g